OpenStudy (anonymous):

Solve 3x^{2}+4x-4=0 by completing the square. Please explain the steps please. I got lost in the part where you're supposed to divide by 2 and square b?

OpenStudy (anonymous):

can someone help @amistre64 @Hero @ash2326 @experimentX

hartnn (hartnn):

first make the co-efficient of x^2 as 1 by dividing both sides by 3, what u get ?

OpenStudy (anonymous):



hartnn (hartnn):

correct, so now your co-efficient of x = b = 4/3, divide it by 2, what u get ?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

I don't understand what I have to do for the rest of the problem @hartnn

hartnn (hartnn):

now square that, what u get ?

OpenStudy (anonymous):

you're not supposed to square root? uh, 4/9

hartnn (hartnn):

no, not square root, u need to square it, and 4/9 is correct! so , now add 4/9 on both sides, what u get ?

OpenStudy (anonymous):

x^2 + 4/3x+4/9 = -8/9

hartnn (hartnn):

correct . now u notice that left side is a perfect square and can be written in the form (x+a)^2 can u write it in that form ?

OpenStudy (anonymous):

I need to go, I'll check back later :(

hartnn (hartnn):

no problem.

OpenStudy (sirm3d):

is this problem an unfinished business?

OpenStudy (anonymous):

I don't understand, so (x+x^2)? @sirm3d yeah, I'm still confused on how to solve the rest of the problem :(

OpenStudy (sirm3d):

the left side of the equation should now be \[\left( x+\frac{ 2 }{ 3 } \right)^{2}\]

OpenStudy (sirm3d):

the trick is to get the square root of the first and last terms of the left side of the equation enclose by a paor of parenthesis, and square it

OpenStudy (anonymous):

so I set that left side you gave me and the right to = -8/9

OpenStudy (anonymous):

It's confusing because I had to turn the 2/3^2 to a 4/9 before and now I'm changing it back? @sirm3d

OpenStudy (sirm3d):

it is necessary in completing squares

OpenStudy (sirm3d):

you have to square (2/3) to get a perfect square trinomial expression on the left side of your equation

OpenStudy (anonymous):

\[(x+2/3)^2=-8/9\]

OpenStudy (anonymous):

it would be an i number for the square root of -8/9?

OpenStudy (sirm3d):

yes

OpenStudy (anonymous):

\[x+2/3=i 0.942809042

OpenStudy (anonymous):

\[x+2/3=i0.942809042\]

OpenStudy (anonymous):

that doesn't look right..

OpenStudy (sirm3d):

\[x+\frac{ 2 }{ 3 }=\pm \frac{ 2i \sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

x_x any little short explanation on how you got the right side?

OpenStudy (anonymous):

you're literally saving my life right now lol (I have a quiz tomorrow morning, well today actually cause it's 4am and I'm still awake from a trippy halloween)

OpenStudy (sirm3d):

the square root of a fraction is a fraction of square roots \[\sqrt{\frac{ 8 }{ 9 }}=\frac{ \sqrt{8} }{ \sqrt{9} }\]

OpenStudy (sirm3d):

the i number is necessary because the right side is negative, -8/9

OpenStudy (anonymous):

Oh! so you do each separate not as a whole.

OpenStudy (anonymous):

I could just leave that as my final answer since both are gonna be imaginary right?

OpenStudy (sirm3d):

yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (sirm3d):

it would be best to direct that question to your teacher, since he/she will be the one to check answers